Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 308     Entry time: 17 Aug 2023, 11:07
Author: Teguh Citra Asmara, Eugenio Paris 
Group: SCS 
Subject: Morning Shift Summary on 16.08.2023 

Achievements:

  1. Performed tr-RIXS on peak A, B, and C at 2.4 mJ/cm^2
  2. Performed tr-XAS on peak A at 3 fluences
  3. Performed tr-RIXS on peak A at 4 mJ/cm^2

Problems:

  1. BAM delay line compensator was not working due to incorrect macro. DOC and LM fixed the issue.
  2. Delay line scan of hRIXS camera was not working as it was still configured for Viking. MT fixed the issue.
  3. The new centroiding algorithm (provided natively by the 4137 kernel) is not consistent with integration mode. For the time being, we use an older, manually configured centroiding algorithm developed for a previous user beamtime.
  4. tR-XAS results show large changes, but these large changes were not reflected in tr-RIXS.
ELOG V3.1.4-7c3fd00